Tyler parents pay him $0.50 to do laundry and $1.25 to mow the lawn. In one month, he does the laundry 6 more times than he mows the lawn. If his parents pay him $13.50 that month, how many times did he mow the lawn?
Using guess and check, I got 6 times mowing the lawn for $7.5 and 12 times doing the laundry for $6.00.
How would I set this up without using guess and check?

Answer by richwmiller(17219)

l=laundry
m=lawn
l=6+m
50l+125m=1350
50*(6+m)+125m=1350
300+50m+125m=1350
175m=1050
m=6
l=6+6=12